OpenStudy (anonymous):

What are the solutions of x2 + 8x − 20 = 0? Show step

OpenStudy (anonymous):

you can factorize it or use quadratic formula

OpenStudy (anonymous):

first i say the solution x=2

OpenStudy (anonymous):

i hjave the answer but i need the steps

OpenStudy (anonymous):

x2+8x-20=0 10x=0+20 x=20/10 =2

OpenStudy (anonymous):

actually the answer is x+ -10, x= 2

OpenStudy (anonymous):

*x= -10

OpenStudy (anonymous):

This formula can solve any quadratic equation \[\LARGE \frac{ -b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

in the question -20

OpenStudy (anonymous):

any @kiamonstaa u get answers steps

OpenStudy (anonymous):

sorry anyway

OpenStudy (anonymous):

@thushananth01 so whats the c?

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

\[\large ax^{2}+bx +c\]

OpenStudy (anonymous):

oh okay thanks =)

OpenStudy (anonymous):

np